OpenStudy (anonymous):

Prove that a line parallel to one side of a triangle divides the other two sides proportionally. Be sure to create and name the appropriate geometric figures

OpenStudy (anonymous):

I have no clue :(

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Given triangle ABC with segment DE parallel to side B.

OpenStudy (mathstudent55):



OpenStudy (anonymous):

I don't get why the numbers are there if we just need to prove that the two sides are proportional @mathstudent55

OpenStudy (mathstudent55):

Using corresponding angles and the parallel lines, angles 1 and 2 are congruent. Also, angles 3 and4 are congruent. By AA Similarity, triangles ABC and ADE are similar. That makes the lengths of corresponding sides proportional. \(\dfrac{AD}{AB} = \dfrac{AE}{AC} \) By the segment addition postulate, we can rewrite segments AB and AC as sums: \(\dfrac{AD}{AD + DB} = \dfrac{AE}{AE + EC} \)

OpenStudy (mathstudent55):

I numbered the angles to make it easier to mention them instead of using three-letter names. Angle 1 is easier to write than angle ADE.

OpenStudy (anonymous):

ohh ok I see

OpenStudy (mathstudent55):

We can now take the reciprocal of both sides: \(\dfrac{AD + DB}{AD} = \dfrac{AE + EC}{AE} \) By a property of proportions, we can conclude: \(\dfrac{DB}{AD} = \dfrac{EC}{AE} \)

OpenStudy (mathstudent55):

That proves it.

OpenStudy (anonymous):

thank you so much .. understand it a lot better now :)

OpenStudy (mathstudent55):

You are welcome.